# Forecasting Melting Points in Svalbard, Norway Using Quantile Gradient Boosting and Adaptive Conformal Prediction Regions

### v1

@@ Abstract
Using data from the Longyearbyen weather station, quantile gradient boosting (“small AI”) is applied to forecast daily 2023 temperatures in Svalbard, Norway. The 0.60 quantile loss weights underestimates about 1.5 times more than overestimates. Predictors include five routinely collected indicators of weather conditions, each lagged by 14 days, yielding temperature forecasts with a two-week lead time. Conformal prediction regions quantify forecasting uncertainty with provably valid coverage. Forecast accuracy is evaluated with attention to local stakeholder concerns, and implications for Arctic adaptation policy are discussed.
conformal prediction regions,
@@ 1 Introduction
A key feature of the IPCC report is its reliance on forecasts at very large spatial and temporal scales, consistent with planet-wide coverage and the gradual pace of climate change. Spatial scales can span thousands of square kilometers. [^1] For example, the Community Earth System Model version 2 (CESM2) is typically run at nominal $1^{\circ}$ resolution, corresponding to grid cells of roughly 10,000–15,000 km2, depending on latitude. Downscaling methods have been proposed for regional grids, but they rest on strong assumptions about the parent Global Climate Model (GCM), and no consensus has emerged on a preferred approach (Nishizawa et al., 2018). Moreover, finer spatial grids would greatly increase computational demands that are already near prohibitive. The temporal scale reflects grid cells updated internally every 30 minutes, with atmosphere–land–ocean–ice coupling on hourly intervals and outputs usually produced as daily or monthly means (Danabasoglu et al., 2020; National Center for Atmospheric Research, 2020). Projections typically extend years into the future.
@@
The goals of numerical weather prediction (NWP) differ from those of climate modeling and require much finer spatial and temporal detail. [^2] Global weather prediction systems, such as ECMWF’s Integrated Forecasting System (IFS) and NOAA’s Global Forecast System (GFS), use grid cells covering on the order of 80–200 km2. Even the most advanced regional models, such as NOAA’s High-Resolution Rapid Refresh (HRRR) or the Weather Research and Forecasting (WRF) model, employ grid spacings of 1–3 km and still aggregate conditions across areas of 1–9 km2. Their time steps are on the order of seconds to minutes, with outputs every 1–3 hours for global models and hourly (or finer) for regional models (European Centre for Medium-Range Weather Forecasts, 2023; NOAA National Centers for Environmental Prediction, 2023; NOAA Earth System Research Laboratory, 2023). Although NWP calculations are performed at much smaller scales, the resulting forecasts still aggregate across those spatial and temporal units. Forecasts for specific sites represent averages across larger areas, and hourly predictions are aggregates over seconds and minutes.
@@
The coarse spatial and temporal resolution characterizing GCMs and NWP is especially consequential for forecasts in high-latitude regions such as the Arctic, where small temperature differences around the local freezing point can have major implications for phenomena such as “Arctic amplification,” discussed in more detail below. This motivates the development of complementary forecasting approaches that can resolve finer spatial and temporal detail, particularly when informing short-term adaptation measures in specific locales.
@@
There are further difficulties. The conditions forecast are usually standardized. The snowfall expected in, say, Fairbanks, Alaska, “during the next 24 hours” is an example. Different locations, however, may wish to supplement such forecasts. In the Arctic, local officials might want to know well in advance the probability that $0^{\circ}\mathrm{C}$ will be exceeded. Such information could guide the timing of sediment removal, pre-emptive slope closures, and preparation of wastewater systems for meltwater discharge, among other actions.
@@
Estimating forecast uncertainty remains problematic (Fu, 2025). Gettleman and Rood (2016, 12) summarize the challenges for climate models, which apply equally to weather forecasting: “Uncertainty in climate models has several components. They are related to the model itself, to the initial conditions of the model, and to the inputs that affect the model. All three must be addressed for the model to be useful.”
@@
There have been recent technical advances, to be sure (Pathak et al., 2022; Price et al., 2024; Bodnar et al., 2025). Yet issues of coarse spatial scales, standardized output, and uncertain reliability remain largely unresolved (Ortega et al., 2022; Balaji et al., 2022; Chang et al., 2023). This paper uses Svalbard, Norway, as a test case to advance Arctic temperature forecasting while addressing these three limitations. Temperatures are the focus because of growing global concern over rising heat.
@@
Section 2 provides background on Svalbard and the forecasting challenges posed by Arctic amplification. Section 3 describes the data, the construction of holdout samples, and the 14-day lagging of predictors used to support legitimate forecasting. Section 4 introduces the statistical methods, focusing on quantile supervised learning. Section 5 presents the empirical results, emphasizing visualization and model interpretation. Section 6 addresses uncertainty through conformal prediction regions for multiple time-series data and offers a grounded way to communicate forecast reliability to stakeholders. Section 7 discusses policy implications, with emphasis on short-fuse adaptations to Arctic melting, and Section 8 concludes. Pseudocode for the conformal inference procedure appears in the Appendix.
@@ 2 Svalbard, Norway: The Forecasting Setting
The maritime climate is distinctive in part because Svalbard is warmed by Atlantic Ocean currents. Average winter temperatures in Longyearbyen range from about $20^{\circ}\mathrm{C}$ to $14^{\circ}\mathrm{C}$, while summers are cool, typically between $3^{\circ}\mathrm{C}$ and $7^{\circ}\mathrm{C}$. In recent years, temperatures have more often spiked above $10^{\circ}\mathrm{C}$, a trend climate scientists find troubling.
Snow falls during much of the year, but total precipitation is relatively low, technically making Svalbard a polar desert. From late October to mid-February the archipelago experiences the polar night, when the sun does not rise above the horizon. From mid-April to late August the sun remains above the horizon, bringing continuous daylight. In recent decades Svalbard has warmed at a rate several times the global average, with mean winter temperatures increasing by more than $5^{\circ}\mathrm{C}$ since the 1970s. Glaciers are retreating, sea ice is thinning, and thawing permafrost is reshaping the terrain. The archipelago is now among the world’s most affected regions of climate change (Urbański and Litwicka, 2022; Karlsen et al., 2024; Bradley et al., 2025; Schuler et al., 2025).
@@ 2.1 Some Implications for Forecasting Temperatures
A practical attraction of Svalbard for temperature forecasting is the weather station at Longyearbyen Airport, whose data are curated by NOAA in collaboration with its international partners. These data are freely available and easy to download. Under the Svalbard Treaty of 1920, Norway retains sovereignty but grants citizens of other signatory nations equal rights to engage in scientific and commercial activities. Several international research stations are operated by Russia, Poland, Germany, China, the United States, and others.
@@
The same Planck effect also operates higher in the atmosphere. The “effective emission height” is the average altitude from which the planet’s infrared radiation escapes directly to space. As greenhouse-gas concentrations rise, this emission level shifts upward. Because temperature generally decreases with altitude, radiation then originates from colder air. By the same $T^{3}$ relationship, the efficiency of infrared emission is reduced. Thus, the Planck effect contributes to polar amplification in two ways: directly at the cold surface and indirectly aloft through the upward shift of the emission height to colder temperatures. Arctic amplification provides an essential context for the forecasting that follows.
@@ 3 Data
The data come from the Longyearbyen weather station. Observations for 2023 are used for training. Data from 2022 provide “honest” assessments of the number of algorithm iterations required, and the 2024 data serve as a fully pristine holdout sample used as “new cases” to document true forecasting skill. They play no role in training. The assumptions required to combine all three datasets are discussed shortly. Data from 2021 are also used for variable construction, as described below.
@@
The downloaded data provide hourly observations within each day. Hourly data are too detailed for the analyses to follow and add unnecessary complexity; daily data are used instead. The response variable is the daily solar-time 2 p.m. air temperature in degrees Celsius at the Longyearbyen weather station. Daily temperature values at 2 p.m. solar time are taken from the Integrated Surface Database using the worldmet package in R. 2 p.m. value is an instantaneous observation—a snapshot of conditions at a constant daily reference time rather than an hourly or daily mean. The 2 p.m. solar-time convention is used uniformly throughout the year, including during the polar night, to provide a consistent temporal reference for forecasting. This differs from global climate and numerical weather prediction procedures, which typically rely on spatially and temporally averaged fields.
@@
Predictors are all lagged by 14 days. They include (1) wind direction in degrees from true north, (2) wind speed in meters per second, (3) air temperature in degrees Celsius, (4) atmospheric pressure in hectopascals (hPa), (5) visibility in meters, (6) dew point in degrees Celsius, (7) relative humidity in percent, and (8) a day counter ranging from 1 to 365. The counter can capture temporal trends: on average, the diurnal months are warmer than the nocturnal months, although temperature changes over time may be nonlinear. Several predictors are likely to interact in complex ways (Semenov, 2021). The 14-day lags translate into fitted 2 p.m. temperatures two weeks later, yielding forecasts two weeks in advance. [^3] Lagging variables is a routine procedure in feature construction. In this case, however, lagging the first 14 days of a year pushes their reference points into the final 14 days of the previous year. Data from 2021 are therefore included to provide lagged values for the first 14 days of 2022. Otherwise, the 2021 data are not used.
@@
The complete dataset forms a multiple time series, making temporal dependence a potential complication. Holdout samples drawn at random from the 2023 training data would obscure any such dependence (Hyndman and Athanasopoulos, 2021, sec. 5.8). Instead, separate holdouts from 2022 and 2024 are used. The same physical processes should apply during the identical months of 2022, 2023, and 2024; data for all three years are treated as random realizations from the same underlying joint probability distribution. Nonetheless, Arctic amplification may affect this comparability, an issue examined empirically in Section 5.2.
@@ 4.1 Background
Models and algorithms from statistics and computer science are increasingly used in climate and weather applications (Ma et al., 2023; Miloshevich et al., 2024; Kvånum et al., 2025). Especially relevant to the analyses that follow is the work of Velthoen et al. (2023), who analyzed several years of daily precipitation observations from the Dutch KNMI weather-station network. The precipitation distribution exhibits a long right tail. Weather-station data provide one source of predictors; deterministic precipitation forecasts from the ECMWF weather model provide another. Combined, these data form a multiple time series.
The authors developed gbex, a quantile gradient-boosting algorithm (Friedman, 2001, 2002) tailored to rare and extreme precipitation events using conditional quantiles at very high levels (e.g., 0.95 or 0.99). Quantile regression methods typically struggle when either tail is sparse. The gbex procedure borrows strength from Extreme-Value Theory (EVT) by fitting a Generalized Pareto Distribution (GPD) to values exceeding a high threshold while allowing the GPD parameters to depend on predictors via quantile gradient boosting. The approach was “validated” by comparing the gbex results with classical methods such as quantile linear regression (Koenker, 2005), showing superior calibration for the extremes.
There is much to admire in Velthoen et al. (2023), but the analyses that follow depart in several important respects. The setting is above the Arctic Circle, where climate and weather processes differ substantially. Consequently, extreme events are not the primary focus; quantile gradient boosting is employed for other purposes.
No enhancements from parametric statistical procedures are used. As Breiman (2001) emphasized, there is considerable skepticism toward statistical modeling unless it has been thoroughly vetted. Nor is algorithmic tuning employed, to avoid post–model-selection bias (Kuchibhotla, Kolassa and Kuffner, 2022). The goal is to perform genuine forecasting, evaluated with independent holdout samples.
Time-series data often exhibit substantial temporal dependence. To obtain valid assessments of forecasting accuracy and uncertainty, the training procedure is expanded to account for this dependence. The analyses adopt standard time-series methods (Box et al., 2015) as a training enhancement.
Finally, provably valid estimates of forecasting uncertainty are obtained using adaptive conformal prediction regions (Romano, Patterson and Candès, 2019). The requisite exchangeability is addressed explicitly, and supplemental results are provided that may be of particular interest to stakeholders.
4.2 The Statistical Procedures Used

### v2

# Forecasting Arctic Temperatures with Temporally Dependent Data Using Quantile Gradient Boosting and Adaptive Conformal Prediction Regions

@@ Abstract
Using data from the Longyearbyen weather station, quantile gradient boosting (“small AI”) is applied to forecast daily temperatures in Svalbard, Norway. Temperatures above $0^{\circ}\mathrm{C}$ are of special interest because of their impact on ice, snow, and tundra permafrost. To improve forecasting skill for warmer temperatures, the target quantile is 0.60; forecast underestimates are weighted 1.5 times more heavily than forecast overestimates when the quantile loss is computed. Predictors include eight routinely collected indicators of weather conditions, each lagged by 14 days, yielding temperature forecasts with a two-week lead time. Adaptive conformal prediction regions quantify forecasting uncertainty with provably valid coverage. Using a holdout sample, a forecast of $0^{\circ}\mathrm{C}$ is correct 14 days later at least 80% of the time. Implications for Arctic adaptation policy are discussed.
adaptive conformal prediction regions,
@@ 1 Introduction
A key feature of the IPCC report is its reliance on forecasts at very large spatial and temporal scales, consistent with planet-wide coverage and the gradual pace of climate change. Spatial scales can span thousands of square kilometers. [^1] For example, the Community Earth System Model version 2 (CESM2) is typically run at nominal $1^{\circ}$ resolution, corresponding to grid cells of roughly 10,000–15,000 km2, depending on latitude. Downscaling methods have been proposed for regional grids, but they rest on strong assumptions about the parent Global Climate Model (GCM), and no consensus has emerged on a preferred approach (Nishizawa et al., 2018). Moreover, finer spatial grids would greatly increase computational demands that are already near prohibitive. The temporal scale reflects grid cells updated internally every 30 minutes, with atmosphere–land–ocean–ice coupling on hourly intervals and outputs are usually produced as daily or monthly means (Danabasoglu et al., 2020; National Center for Atmospheric Research, 2020). Projections typically extend years into the future.
@@
The goals of numerical weather prediction (NWP) differ from those of climate modeling and provide output at much finer spatial and temporal detail. [^2] Global weather prediction systems, such as ECMWF’s Integrated Forecasting System (IFS) and NOAA’s Global Forecast System (GFS), use grid cells covering on the order of 80–200 km2. Even the most advanced regional models, such as NOAA’s High-Resolution Rapid Refresh (HRRR) or the Weather Research and Forecasting (WRF) model, employ grid spacings of 1–3 km and still aggregate conditions across areas of 1–9 km2. Their time steps are on the order of seconds to minutes, with outputs every 1–3 hours for global models and hourly (or finer) for regional models (European Centre for Medium-Range Weather Forecasts, 2023; NOAA National Centers for Environmental Prediction, 2023; NOAA Earth System Research Laboratory, 2023). Although NWP calculations are performed at much smaller scales, the resulting forecasts still aggregate across those spatial and temporal units. Site-specific forecasts represent averages across larger areas, and hourly predictions are aggregates over seconds and minutes.
@@
The coarse spatial and temporal resolution characterizing GCMs and NWP is especially consequential for forecasts in high-latitude regions such as the Arctic. Small temperature differences around the local freezing point can have important implications, intensified by “Arctic amplification.” Arctic amplification is discussed in more detail shortly.
@@
In addition, the conditions forecasted are usually standardized. The snowfall expected in, say, Fairbanks, Alaska “during the next 24 hours” is one example. Different locations, however, may wish to supplement such forecasts. In the Arctic, local officials might want to know well in advance the probability that $0^{\circ}\mathrm{C}$ will be exceeded. Such information could guide the timing of sediment removal, preemptive slope closures, and preparation of wastewater systems for meltwater discharge, among other actions.
@@
Estimating forecasting uncertainty is problematic as well (Fu, 2025). Gettleman and Rood (2016, 12) summarize the challenges for climate models, which apply equally to weather forecasting: “Uncertainty in climate models has several components. They are related to the model itself, to the initial conditions of the model, and to the inputs that affect the model. All three must be addressed for the model to be useful.”
@@
There have been recent technical advances (Pathak et al., 2022; Price et al., 2024; Bodnar et al., 2025). Yet issues of coarse spatial scales, standardized output, equivocal reliability, and other issues remain largely unresolved (Ortega et al., 2022; Balaji et al., 2022; Chang et al., 2023; Zhang et al., 2025). This paper uses Svalbard, Norway, as a test case to advance Arctic temperature forecasting while addressing these three limitations. Temperatures are the focus because of legitimate concerns about a warming climate.
The statistical methods used here are best viewed as a complement to techniques usually applied at larger spatial and temporal scales. Because these methods are more surgical, they can resolve finer spatial and temporal detail. One can obtain legitimate forecasts coupled with valid measures of forecasting uncertainty well suited for highly local and rapid adaptations to Arctic warming.
@@
Section 2 provides background on Svalbard and the forecasting challenges posed by Arctic amplification. Section 3 describes the data, the construction of holdout samples, and the 14-day lagging of predictors used to support legitimate forecasting. Section 4 introduces the statistical methods, focusing on quantile supervised learning. Section 5 presents the empirical results, emphasizing visualization and algorithm interpretations. Section 6 addresses uncertainty through adaptive conformal prediction regions for multiple time-series data and offers a grounded way to communicate forecast reliability to stakeholders. Section 7 discusses policy implications, with emphasis on short-fuse adaptations to Arctic melting, and Section 8 concludes. A synopsis of the data analysis steps is provided in Appendix A.
@@ 2 Svalbard, Norway: The Forecasting Setting
The maritime climate is distinctive in part because Svalbard is warmed by Atlantic Ocean currents. Average winter temperatures in Longyearbyen range from about $20^{\circ}\mathrm{C}$ to $14^{\circ}\mathrm{C}$, while summers are typically between $3^{\circ}\mathrm{C}$ and $7^{\circ}\mathrm{C}$. In recent years, summer temperatures have more often spiked above $10^{\circ}\mathrm{C}$, a trend climate scientists find troubling.
Snow falls during much of the year, but total precipitation is relatively low – technically making Svalbard a polar desert. From late October to mid-February the archipelago experiences the polar night, when the sun does not rise above the horizon. From mid-April to late August, the sun remains above the horizon, bringing continuous daylight. In recent decades Svalbard has warmed at a rate several times the global average, with mean temperatures increasing by more than $5^{\circ}\mathrm{C}$ since the 1970s. Glaciers are retreating, sea ice is thinning, and thawing permafrost is reshaping the terrain. The archipelago is now among the world’s regions most affected by climate change (Urbański and Litwicka, 2022; Karlsen et al., 2024; Bradley et al., 2025; Schuler et al., 2025).
@@ 2.1 Some Implications for Forecasting Temperatures
A practical attraction of Svalbard for temperature forecasting is the weather station at Longyearbyen Airport, whose data are curated by NOAA in collaboration with its international partners. These data are freely available and easy to download. Under the Svalbard Treaty of 1920, Norway retains sovereignty but grants citizens of other signatory nations equal rights to engage in scientific and commercial activities. International research stations are operated by Russia, Poland, Germany, China, the United States, and others.
@@
The same Planck effect also operates higher in the atmosphere. The “effective emission height” is the average altitude from which the planet’s infrared radiation escapes directly to space. As greenhouse-gas concentrations rise, this emission level shifts upward. Because temperature generally decreases with altitude, radiation then originates from colder air. By the same $T^{3}$ relationship, the efficiency of infrared emission is reduced. Thus, the Planck effect contributes to polar amplification in two ways: directly at the cold surface and indirectly aloft through the upward shift of the emission height to colder temperatures. Arctic amplification provides a potential complication for the forecasting that follows.
@@ 3 Data
The data come from the Longyearbyen weather station at the airport in Svalbard, Norway, and are easily downloaded from the Integrated Surface Database using the worldmet package in R. The database contains weather station data from around the globe. It uses the same standard format regardless of origin.
Anticipating the use of adaptive conformal prediction regions (Romano, Patterson and Candès, 2019), a variant on split-sample methods is employed. Observations for 2023 constitute the training data. Calibration data from 2022 provide an “honest” assessment of the number of algorithm iterations required. They are also the observations needed to construct proper nonconformal scores.
The calibration data are divided into observations for which the forecasted temperature is $0^{\circ}\mathrm{C}$ and observations for which the forecasted temperature is $\leq 0^{\circ}\mathrm{C}$. The melting of snow, sea ice, glacier ice, and permafrost creates positive feedback loops that change the manner in which temperature variation is produced. On subject-matter and statistical grounds, adaptive conformal methods are therefore applied separately to each subset. [^3] One might wonder why the data splits are made using forecasted temperatures rather than observed temperatures. Observed temperatures are certainly available in the weather station data. But when forecasting, only the forecast is known. If the future temperature were known, there would be no need to forecast it.
Data from 2024 serve as a pristine holdout sample treated as “new cases” to document true forecasting skill. As test data, these observations have no role whatsoever in training. The data split by forecasted temperature applied to the calibration data is applied again. The requirements necessary to capitalize properly on all three datasets within the split-sample approach are discussed in Section 6. Some data from 2021 are used for lagged-variable construction, briefly explained in footnote 4. They play no other role in the analyses.
The training data, calibration data, and test data from the Integrated Surface Database come with hourly observations within each day. Hourly data are too detailed for the analyses to follow and add unnecessary complexity; daily data are used instead.
@@
The response variable is the daily, solar-time 2 p.m. air temperature in degrees Celsius at the Longyearbyen weather station. Each daily temperature value at 2 p.m. solar time is an instantaneous observation—a snapshot of conditions with a constant daily reference time rather than an hourly or daily mean. The 2 p.m. solar-time convention is used uniformly throughout the year, including during the polar night, to provide a consistent temporal reference for forecasting. This differs from global climate and numerical weather prediction procedures, which typically rely on spatially and temporally averaged fields.
@@
Predictors are all lagged by 14 days because that seems to be a lag commonly used by climate scientists (Li et al., 2024). In future work, longer lags could be considered. The predictors include (1) wind direction in degrees from true north, (2) wind speed in meters per second, (3) air temperature in degrees Celsius, (4) atmospheric pressure in hectopascals (hPa), (5) visibility in meters, (6) dew point in degrees Celsius, (7) relative humidity in percent, and (8) a day counter ranging from 1 to 365. The counter can capture temporal trends: on average, the diurnal months are warmer than the nocturnal months, although temperature changes over time may be nonlinear. Several predictors are likely to interact in complex ways (Semenov, 2021). The 14-day lags translate into fitted 2 p.m. temperatures two weeks later that are a foundation for forecasts 14 days in advance. [^4] Lagging variables is a routine procedure in feature construction. In this case, however, lagging the first 14 days of a year pushes their reference points into the final 14 days of the previous year. Data from 2021 are therefore included to provide lagged values for the first 14 days of 2022. Otherwise, the 2021 data are not used.
@@
The complete dataset forms a multiple time series, making temporal dependence a potential complication. Split samples are commonly disjoint subsets chosen at random from the data available, but that is likely to obscure any temporal dependence (Hyndman and Athanasopoulos, 2021, sec. 5.8). A key requirement for the 2022 and 2024 holdout samples is that the same physical processes apply during the identical months of 2022, 2023, and 2024; data for all three years are treated as random realizations from the same underlying joint probability distribution. Nonetheless, Arctic amplification may affect this comparability, an issue examined empirically in Section 5.2.
@@ 4 Statistical Methods
The analyses to follow produce legitimate forecasts from supervised machine learning applied to temporal data. Holdout samples are used to construct provably valid prediction regions. The statistical framework is an interlocking set of procedures. After a search in Google Scholar, their combination appear to be novel. Each component is discussed in a grounded manner when deployed. Appendix A provides a step-by-step synopsis.
@@ 4.1 Background
Machine learning algorithms from statistics and computer science are increasingly used in climate and weather applications, including forecasting (Ma et al., 2023; Miloshevich et al., 2024; Kvånum et al., 2025; Zhang et al., 2025). Especially relevant to the analyses that follow is the work of Velthoen et al. (2023), who analyze several years of daily precipitation observations from the Dutch KNMI weather-station network. The precipitation distribution exhibits a long right tail. Weather-station data provide one source of predictors; deterministic precipitation forecasts from the ECMWF weather model provide another. Combined, these data form a multiple time series.
The authors develop gbex, a quantile gradient-boosting algorithm (Friedman, 2001, 2002) tailored for rare and extreme precipitation events using conditional quantiles at very high levels (e.g., 0.95 or 0.99). Quantile regression methods typically struggle when either tail is sparse. The gbex procedure borrows strength from Extreme-Value Theory (EVT) by fitting a Generalized Pareto Distribution (GPD) to values exceeding a high threshold while allowing the GPD parameters to depend on predictors via quantile gradient boosting. The approach is “validated” by comparing the gbex results with classical methods, such as quantile linear regression (Koenker, 2005), showing superior calibration for the extremes.
There is much to admire in Velthoen et al. (2023), but the analyses to follow depart in several important respects. The setting is above the Arctic Circle, where climate and weather processes differ substantially. Consequently, extreme events are not the primary focus; quantile gradient boosting is employed for other purposes.
No enhancements from parametric statistical procedures are used. As Breiman (2001) emphasizes, there is considerable skepticism about statistical modeling unless it has been thoroughly vetted. Nor is algorithmic tuning employed. Therefore, post–model-selection bias is avoided (Kuchibhotla, Kolassa and Kuffner, 2022).
Time-series data commonly exhibit substantial temporal dependence. To obtain valid assessments of forecasting accuracy and uncertainty, the training procedure is expanded to account for this dependence. The analyses adopt standard time-series methods (Box et al., 2015) as a training extension.
Finally, valid estimates of forecasting uncertainty are obtained using adaptive conformal prediction regions (Romano, Patterson and Candès, 2019). The requisite exchangeability is addressed explicitly. Supplemental results are provided that may be of particular interest to stakeholders.
4.2 Some Details